# The continuous spectrum of bound states in expulsive potentials

### v1

@@ III Bound states in the 1D nonlinear Schrödinger equation (NLSE) with the expulsive potential
While the stability of the existing eigenstates in the framework of the linear Schrödinger equation is obvious, it is a nontrivial problem in the case of the NLSE. Numerical simulations of the perturbed evolution of bound states produced by Eq. (9) with the cubic nonlinearity, $\sigma=1$ and $g=\pm 1$, has demonstrated that all the respective bound states remain stable, while their stationary shape is slightly deformed by the cubic term. A typical example of the stable evolution of the bound state for g=1g=-1 (the self-focusing cubic term), $\gamma=2$ (the quartic expulsive potential), and E=1E=-1 is presented in Fig. 6.
Figure 6: The stable evolution of the spatially even eigenstate obtained as the numerical solution of Eq. (10) with $\gamma=2$, g=1g=-1, $\sigma=1$ (the cubic self-focusing term), and E=1E=-1. The norm of the eigenstate is $N=2.86$.
@@
The 1D spatially odd (dipole) eigenstates also remain stable, being only weakly deformed by the cubic nonlinearity. An example is displayed in Fig. 7 for $\gamma=2$, $\sigma=1$, $g=+1$ (the self-defocusing cubic term) and $E=0$, i.e., the same parameters as those (except for $g=0$) of the dipole-mode eigenstate shown in Fig. 3(b), which was produced by the linear version of Eq. (10). Note that both eigenstates are very well approximated by the asymptotic expression (11) with appropriate values of the fitting constants, viz., $\chi_{0}=\pi/3$ for $g=0$ and $\chi_{0}=5\pi/12$ for $g=+1$.
@@
Figure 7: The stable spatially odd eigenstate produced by the numerical solution of Eq. (10) with $g=+1$, $E=0$, and $\gamma=2$, cf. its counterpart plotted in Fig. 3(b) produced for the same parameters by the linearized equation (with $g=0$). The norm of eigenstate is $N=1.23$. The dashed curve represents the corresponding asymptotic approximation (11) with fitting constants $\chi_{0}=5\pi/12$ and $\varphi_{0}=0.995$.
In the case of the quintic self-focusing, which corresponds to g=1g=-1 and $\sigma=2$ in Eq. (9), the eigenstate whose norm exceeds a certain critical value is unstable against the onset of the collapse. This point will be reported in detail elsewhere.
IV Two-dimensional bound states
@@
The natural 2D version of Eq. (9) can also be realized experimentally as the GPE for ultracold atoms, under the action of the isotropic expulsive optical potential, induced by the blue-detuned laser beam. In terms of the polar coordinates $\left(r,\theta\right)$, the 2D GPE is written as
@@
with $\gamma\geq 1$ (here, only the cubic nonlinear term is considered, if any). The stationary version of Eq. (16) with vorticity $S$ (alias magnetic quantum number) and eigenvalue $E$ (the chemical potential, in terms of the BEC), is produced by the substitution
@@
Similar to Eq. (11) in the 1D case, it is straightforward to construct the asymptotic form of the tail of the solution to Eq. (18) at $r\rightarrow\infty$, in the case of $\gamma>1$. The first two terms of the asymptotic expansion are
@@
cf. its 1D counterpart (12). As above, $\varphi_{0}$ and $\chi_{0}$ are indefinite constants, in the framework of the asymptotic expansion at $r\rightarrow\infty$. Note that only the smallest asymptotic term (20) includes the vorticity.
As seen from expression (19), the 2D norm of these eigenstates, $N_{\mathrm{2D}}=2\pi\int_{0}^{\infty}$ $\varphi^{2}(r)rdr$, converges under precisely the same condition as in 1D, viz., $\gamma>1$, i.e., if the 2D expulsive potential is steeper than the isotropic anti-HO potential. A qualitative explanation for this counter-intuitive conclusion is the same as in the 1D setting: under the action of the expulsive potential, a classical particle would be rolling down the steep potential hill (along a spiral trajectory, if the particle carries the angular momentum, corresponding to $S\geq 1$), with rapid acceleration. In terms of the particle’s quantum counterpart, the acceleration gives rise to the strongly oscillating phase in the wave function, thus leading to the normalizability (effective localization) of the eigenstate.
Similar to the 1D case, the 2D Schrödinger equation (16) with $\gamma=1$, the asymptotic expression (19) is relevant solely for $E=0$, as otherwise the second term diverges. For $E\neq 0$, the relevant asymptotic approximation is
$\varphi_{\mathrm{asympt}}^{\mathrm{(2D)}}(r)=\varphi_{0}r^{-1}\cos\left(\frac{r^{2}}{2}+E\ln\left(\frac{r}{l}\right)\right),$ (21)
where $l$ is the radius of the inner core, cf. Eq. (14). Also similar to its 1D counterpart, in the case of $\gamma=1$ the norm of the 2D eigenstate is slowly diverging, $N_{\mathrm{2D}}\simeq\pi\varphi_{0}^{2}\ln\left(L/l\right)$ (cf. Eq. (13)), where $L$ is the radius of the integration domain.
@@
where $\varphi_{0}$ is an arbitrary amplitude (the correctness of solution (23) was also checked by comparing to its numerically constructed counterpart).
For all integer values $S\geq 2$, the norm of the exact wave function (23) converges, yielding
@@
where $\Gamma\left(1/S\right)$ is the Gamma-function. In agreement with the above conclusion, the norm diverges for $\gamma=1$, i.e., $S=1$, see Eq. (22), the respective exact solution (23) taking the form of
@@
This solution is plotted in Fig. 8.
Figure 8: The absolute value $\left|\varphi(r)\right|$ of the exact solution (25) of Eq. (18) with $\gamma=1$, $g=0,$ $S=1$, and $E=0$.
Lastly, if the cubic term is retained in Eq. (18), one can construct an approximate solution for $E=0$ and $S\neq 1$, neglecting the third harmonic in the elementary formula $\sin^{3}\Phi=(3/4)\sin\Phi-(1/4)\sin\left(3\Phi\right)$. The approximate solution is given by Eq. (25) (even if $S$ is not $1$), in which the amplitude is determined by the balance of the linear and nonlinear terms in Eq. (18):
$\varphi_{0}^{2}=-2\left(3g\right)^{-1}\left(S^{2}-1\right).$ (26)
@@
In the case of g=1g=-1 (self-focusing), it is necessary to test the stability of the nonlinear vortex solutions against spontaneous splitting, cf. Refs. [39, 40, 41, 42]. These results will be reported elsewhere.
V Conclusion
@@
Our analysis demonstrates that, on the contrary to the intuitive expectation, the 1D and 2D Schrödinger equations with the expulsive potential which is steeper than the anti-HO (anti-harmonic-oscillator) one, generates a full continuous spectrum of the normalizable bound states. In 1D these may be fundamental and dipole modes (spatially even and odd ones, respectively; plausibly, higher-order excited states exist too, which is a subject for additional analysis). The 2D bound states may carry any integer vorticity (angular momentum). Universal asymptotic expressions for the 1D and 2D states, valid at $|x|\rightarrow\infty$ and $r\rightarrow\infty$, respectively, are found, providing an extremely accurate approximation to the numerically found wave functions. In addition, in the 2D case special vorticity-carrying solutions are found in the exact form. These results essentially extend the concept of BIC (bound states in the continuum), which has recently drawn much interest in the context of photonics, where the paraxial propagation models amount to the same Schrödinger equations as in quantum mechanics. In the limit of the anti-HO shape of the expulsive potential, the respective normalization integrals are weakly divergent, in 1D and 2D settings alike. The nonlinear version of the 1D and 2D models, in the form of the respective GPEs (Gross-Pitaevskii equations), are considered too. In 1D, the cubic self-focusing or defocusing nonlinearity slightly deforms the linear solutions and does not break their stability. The quintic self-focusing term makes the 1D bound states unstable against the onset of the collapse if the norm of the state exceeds the critical value.
As the further extension of the present work, it may be relevant to develop the analysis of the 2D model with the cubic nonlinear term. In particular, a challenging issue is the (in)stability of the nonlinear vortex eigenstates against spontaneous splitting by the azimuthal modulational instability in the case of the self-focusing sign of the nonlinearity, cf. works [39, 40, 41, 42], in which a similar problem was investigated for nonlinear vortex states trapped in the usual isotropic HO potential. In parallel, and from the dynamical systems perspective, it may be relevant to perform a spectral stability analysis (in the realm of the Bogoliubov-de Gennes analysis) [43, 44] in order to track the states’ spectrum and identify potential bifurcations. Finally, a true challenge while numerically computing such states herein is the need of using a large number of grid points for capturing the tails’ oscillations as we observed in the present study. A potential way out for the 2D problem may be the use of mesh adaptation, see, e.g., [45] that will help resolving these oscillations very well with less computational cost. Those studies are underway and results will appear in future contributions.
VI Funding
@@
EGC acknowledges support from the San Diego State University, Department of Mathematics and Statistics startup fund. EGC thanks P.G. Kevrekidis for fruitful discussions during the early stages of this work.
VII Author Contributions
@@ VIII Conflict of Interests
IX Data Availability Statement

### v2

@@ II.1 One-dimensional linear bound states
While the stability of the existing eigenstates in the framework of the linear Schrödinger equation is obvious, it is a nontrivial problem in the case of the NLSE. The numerical solution of the stationary version of Eq. (9) with the cubic nonlinearity, i.e., $\sigma=1$ and $g=\pm 1$, has demonstrated the existence of the corresponding bound states, with profiles somewhat deformed by the cubic term. Simulations of the perturbed evolution of the bound states have demonstrated that they remain stable in some parameter region, and become unstable in another, as briefly demonstrated below.
A typical example of the stable evolution of the bound state for g=1g=-1 (the self-focusing cubic term), $\gamma=2$ (the quartic expulsive potential), and E=0.9E=-0.9 is presented in Fig. 7(a). In this case, the stationary even eigenstate (with $\varphi(-x)=\varphi(x)$) was obtained, as the solution of the corresponding ordinary differential equation (10), by dint of the Runge-Kutta method with the initial conditions $\varphi(x=0)=1.381$ and $d\varphi/dx(x=0)=0$, cf. Eq. (16). The stationary wave function vanishes at edges of the solution domain, $|x|=L=4.8981$. The respective norm of the eigenstate, $\int_{-L}^{+L}\varphi^{2}(x)dx$ (cf. Eq. (11)) is
$N_{\mathrm{1D}}=1.73.$ (19)
To confirm the stability of this eigenstate, we have performed simulations of its perturbed evolution in the framework of the full equation (9), using the split-step Fourier method in the domain $|x|\leq L$, with zero boundary conditions, $\psi(x=\pm L)=0$. The number of Fourier modes was $8192$, and the marching timestep was $\Delta t=10^{-5}$. As the local wavenumber of the oscillating tail increases with $|x|$, more Fourier modes are necessary for the simulations of a larger system. Here, we present the results of the simulations performed in the domain of size $2L=9.7962$, which maintain sufficiently high numerical accuracy. As the initial condition, we took
$\psi(x,t=0)=\varphi(x)+0.001\sin(2\pi x/L),$ (20)
where the second term introduces a small spatially odd perturbation, which breaks the parity of the nonlinear even eigenstate $\varphi(x)$. The aim was, in particular, to test the stability of the eigenstate against the spatial-symmetry-breaking perturbations. Eight snapshots of instantaneous profiles $|\psi(x,t)|$, produced by the simulations at $t=0,2,\cdots,14$, are plotted in Fig. 7(a). The presented results clearly confirm the stability of the nonlinear eigenstate.
A symmetry-breaking instability of the eigenstate occurs for larger absolute values of eigenvalue $E$, An example is presented in Fig. 7(b) for g=1g=-1 and E=1.1E=-$1.$ The corresponding stationary eigenstate is produced by the numerical solution of Eq. (10) with the initial conditions $\varphi(x=0)=1.381$ and $d\varphi/dx\left(x=0\right)=0$, in the domain $0<x<L=4.9253$. The norm of the eigenstate is $N_{\mathrm{1D}}=3.07$, cf. the above value (19). Then, the perturbed evolution of the eigenstate was simulated with the input taken as per Eq. (20) (as above). Figure  7(b) shows snapshots of $|\psi(x,t)|$ at $t=6,8,\cdots,20$. It is seen that, unlike the stable evolution displayed in Fig. 7(a), the odd perturbation grows, so that the central-peak’s position moves to the right and then returns to the origin, due to the interaction with the the oscillatory tail.
The motion of the center of mass of the solutions,
$\langle x\rangle=\frac{1}{N_{\mathrm{1D}}}\int_{-L}^{+L}x|\psi(x)|^{2}dx,$ (21)
is plotted in Fig. 7(c) for the above-mentioned perturbed eigenstates with E=0.9E=-0.9 (the dotted red line), E=1.1E=-1.1 (the dashed blue line), and also for the intermediate case, with E=1.0E=-1.0, with $L=4.9116$ and $N_{\mathrm{1D}}=2.86$ (the solid green line). It is concluded that the parity-breaking instability of the spatially even nonlinear eigenstates occurs at $E\approx-1$, a natural conclusion being that the self-focusing nonlinearity leads to the destabilization of the bound states above a critical value of the norm. In particular, the dashed blue trajectory in Fig. 7(c) demonstrates that cycle of the instability development, which is displayed in Fig. 7(b), repeats with approximate periodicity. Detailed studies of the stability of the nonlinear eigenstates, and production of the respective stability charts for the systems with of the cubic and quintic nonlinear terms, will be a subject of another work.
Figure 7: (a) The stable evolution of the spatially even eigenstate obtained as the numerical solution of Eq. (10) with $\gamma=2$ , g=1g=-1, $\sigma=1$ (the cubic self-focusing nonlinearity), eigenvalue E=0.9E=-0.9, and norm $N_{\mathrm{1D}}=1.73$. Eight snapshots of $|\psi(x,t)|$ at $t=0,2,\cdots,14$ are plotted. (b) Snapshots of $|\psi\left(x,t\right)|$ at $t=6,8,\cdots,20$ illustrate the onset of the spontaneous parity-breaking instability of the even eigenstate with E=1.1E=-1.1 and $N_{\mathrm{1D}}=3.07$. (c) The center-of-mass coordinate (21) as a function of. time for the perturbed evolution of the eigenstates with E=0.9E=-0.9 (the dotted red line), E=1E=-1 (the solid green line), and E=1.1E=-1.1 (the dashed blue line).
@@
The 1D spatially odd (dipole) eigenstates are also weakly deformed by the cubic nonlinearity, and may remain stable, depending on the parameters. An example is displayed in Fig. 8 for $\gamma=2$, $\sigma=1$, $g=+1$ (the self-defocusing cubic term) and $E=0$, i.e., the same parameters as those (except for $g=0$) of the dipole-mode eigenstate shown in Fig. 4(b), which was produced by the linear version of Eq. (10). Note that both eigenstates are very well approximated by the asymptotic expression (12) with appropriate values of the fitting constants, viz., $\chi_{0}=\pi/3$ for $g=0$ and $\chi_{0}=5\pi/12$ for $g=+1$.
@@
Figure 8: The stable spatially odd eigenstate produced by the numerical solution of Eq. (10) with $g=+1$, $E=0$, and $\gamma=2$, cf. its counterpart plotted in Fig. 4(b) produced for the same parameters by the linearized equation (with $g=0$). The norm of the eigenstate is $N=1.23$. The dashed curve represents the corresponding asymptotic approximation (12) with fitting constants $\chi_{0}=5\pi/12$ and $\varphi_{0}=0.995$.
II.3 Two-dimensional bound states
@@
The natural 2D version of Eq. (9) can also be realized experimentally as the GPE for ultracold atoms, under the action of the isotropic expulsive optical potential, induced by the blue-detuned laser beam, or the red-detuned one with intrinsic vorticity. In terms of the polar coordinates $\left(r,\theta\right)$, the 2D GPE is written as
@@
with $\gamma\geq 1$ (here, only the cubic nonlinear term is considered, if any). The stationary version of Eq. (22) with integer vorticity $S$ (alias magnetic quantum number) and eigenvalue $E$ (the chemical potential, in terms of the BEC), is produced by the substitution
@@
Similar to Eq. (12) in the 1D case, it is straightforward to construct the asymptotic form of the tail of the solution to Eq. (24) at $r\rightarrow\infty$, in the case of $\gamma\neq 1$. The first two terms of the asymptotic expansion are
@@
cf. its 1D counterpart (13). As above, $\varphi_{0}$ and $\chi_{0}$ are indefinite constants, in the framework of the asymptotic expansion at $r\rightarrow\infty$. Note that only the higher-order correction to the asymptotic expansion, given by Eq. (26), includes the vorticity.
As seen from expression (25), the 2D norm of these eigenstates, $N_{\mathrm{2D}}=2\pi\int_{0}^{\infty}$ $\varphi^{2}(r)rdr$, converges under precisely the same condition as in 1D, viz., $\gamma>1$, i.e., if the 2D expulsive potential is steeper than the isotropic anti-HO (quartic) potential. A qualitative explanation for this counter-intuitive conclusion is the same as in the 1D setting: under the action of the expulsive potential, a classical particle would be rolling down the steep potential hill (along a spiral trajectory, if the particle carries the angular momentum, corresponding to $S\geq 1$ in expression (23)), with rapid acceleration. In terms of the particle’s quantum counterpart, the acceleration gives rise to the strongly oscillating phase of the wave function, thus leading to the effective linear self-trapping (localization) of the eigenstate, with the normalizable wavefunction.
Similar to the 1D case, for the 2D Schrödinger equation (22) with $\gamma=1$ (the anti-HO expulsive potential), the asymptotic expression (25) takes a different form, with the logarithmic correction to the phase of the rapid oscillations:
$\varphi_{\mathrm{asympt}}^{\mathrm{(2D)}}(r;\gamma=1)=\varphi_{0}r^{-1}\cos\left(\frac{r^{2}}{2}+E\ln\left(\frac{r}{l}\right)\right),$ (27)
where $l$ is the radius of the inner core, cf. Eq. (14). Also similar to its 1D counterpart, in the case of $\gamma=1$ the norm of the 2D eigenstate is weakly divergent, $N_{\mathrm{2D}}\simeq\pi\varphi_{0}^{2}\ln\left(L/l\right)$ (cf. Eq. (15)), where $L$ is the radius of the integration domain.
@@
where $\varphi_{0}$ is an arbitrary amplitude. For all values $S\geq 2$, the norm of the exact wave function (29) converges, yielding
@@
where $\Gamma\left(1/S\right)$ is the Gamma-function. In agreement with the above conclusions, the norm diverges for $\gamma=1$, i.e., $S=1$, see Eq. (28), the respective exact solution (29) taking the form of
@@
This analytical solution and the corresponding numerical solution of Eq. (24), with $\varphi_{0}=0.409$, are plotted by the solid and dashed lines, respectively, in Fig. 9(a), where they completely overlap (actually, the coincidence with the exact analytical solution corroborates the accuracy of the numerical solution). The 2D structure of this vortex eigenstate is illustrated by means of the color plot in Fig. 9(b).
[…]
Lastly, if the cubic term is retained in Eq. (24) with $\gamma=1$ (the anti-HO expulsive potential), one can construct an approximate solution for $E=0$ and $S\neq 1$, neglecting the third harmonic in the elementary formula $\sin^{3}\Phi=(3/4)\sin\Phi-(1/4)\sin\left(3\Phi\right)$. The respective approximate solution is given by Eq. (31) (even if $S$ is not $1$), in which the amplitude is determined by the balance of the linear and nonlinear terms in Eq. (24):
$\varphi_{0}^{2}=2\left(3g\right)^{-1}\left(1-S^{2}\right).$ (32)
@@
In the case of g=1g=-1 (self-focusing), it is necessary to test the stability of the nonlinear vortex solutions against spontaneous splitting, cf. Refs. [57, 58, 59]. These results will be reported elsewhere.
III Conclusions
@@
Our analysis demonstrates that, on the contrary to the intuitive expectation, the 1D and 2D Schrödinger equations with the expulsive potential which is steeper than the quadratic (anti-HO) one, generates a full continuous spectrum of the normalizable bound states, which may be considered as a manifestation of the effective self-trapping in the linear system. In 1D, these bound states may be the fundamental and dipole ones (spatially even and odd modes, respectively; plausibly, higher-order excited states exist too, which is a subject for additional analysis). The 2D bound states may carry any integer vorticity (angular momentum). Universal asymptotic expressions for the 1D and 2D states, valid at $|x|\rightarrow\infty$ and $r\rightarrow\infty$, respectively, are found, providing an extremely accurate approximation to the numerically found wave functions. In addition, in the 2D case special vorticity-carrying solutions are found in the exact form. These results essentially extend the concept of BIC (bound states in the continuum), which has recently drawn much interest in the context of photonics, where the paraxial-propagation models amount to the same Schrödinger equations as in quantum mechanics. In the case of the anti-HO (quadratic) shape of the expulsive potential, the normalization integrals for the bound states are weakly divergent, in 1D and 2D settings alike. The nonlinear version of the 1D and 2D models, in the form of the respective GPEs (Gross-Pitaevskii equations), are briefly considered too. In 1D, the cubic self-focusing or defocusing nonlinearity slightly deforms the linear solutions. A parity-breaking instability occurs in the case of the self-focusing nonlinearity if the norm of the spatially even eigenstate exceeds a critical value. A detailed analysis of the nonlinear models will be reported elsewhere. In particular, a challenging issue is the (in)stability of 2D nonlinear vortex eigenstates against splitting by the azimuthal modulational instability in the case of the self-focusing cubic nonlinearity. In parallel, it is relevant to perform the spectral stability analysis of the nonlinear eigenstates, in the framework of the Bogoliubov-de Gennes equations for small perturbations [60, 61]. As concerns the development of the numerical analysis of the 1D and 2D eigenstates reported in this work, a challenging issue is the need of using a large number of grid points for capturing rapid oscillations of the decaying tails. A potential way to improve this aspect of the 2D solutions may be the use of the mesh adaptation [62], that should help resolving the oscillations very accurately with less computational cost. These studies are currently underway and will be reported elsewhere.
IV Acknowledgments
We thank P. G. Kevrekidis and D. H. J. O’Dell for helpful discussions.
V Funding
@@
E.G.C. acknowledges support from the San Diego State University, Department of Mathematics and Statistics startup fund.
VI Author Contributions
@@ VII Conflict of Interests
VIII Data Availability Statement